OpenStudy (georgia412):

Find the equation of the straight line when gradient =2/5 and passes through point (0,-1)

OpenStudy (zzr0ck3r):

ok sort of like last time, but now they do not give us the y intercept we know the gradient so y = (2/5)x+b now we use the point they gave to find b (0,-1) so when x = 0, y = -1 so y=(2/5)x+b becomes -1 = (2/5)(0)+b what is b?

OpenStudy (georgia412):

Is b 0/5?

OpenStudy (zzr0ck3r):

\[-1=\frac{2}{5}*0+b\] what is b?

OpenStudy (zzr0ck3r):

0*(2/5) = 0

OpenStudy (georgia412):

I see now :) thanks!

OpenStudy (zzr0ck3r):

word \[y=\frac{2}{5}x-1\]